OpenStudy (anonymous):

5900 is invested, part at 10% and part at 8% totaled yeild$534.00. how much was invested at each rate?

OpenStudy (radar):

Lets review what we do know. A total of 34 were sold, the number of white ones plus the number of yellow one equal 34 We know the cost of a sweater if it is white is $9.95, yellow $11.50

OpenStudy (radar):

What are we to find? We want to find the number of each sweater sold. What else do we know? We know that a total $366.20 was sold (white & yellow)

OpenStudy (radar):

Let x = the number of white sweatshirts sold. Taking what we know, we can say 34-x is the number of yellow sold. Does that make sense?

OpenStudy (radar):

It does, so continuing. We now get to the money. If we know the amount of money made for each color, we can get the desired answer.

OpenStudy (radar):

Total money was $366.20 and that would also equal: 9.95x + 11.50 (34-x) = $366.20 Now it is just simple algebra to solve for x (white) and 34-x the yellow ones. Do you need further help?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Have at it, and good luck, post your answer and I will standby.

OpenStudy (anonymous):

i am not good in algebra confused

OpenStudy (radar):

9.95x + 11.50(34) - 11.5x = $366.20 (after clearing the parenthesis -1.55x +391 = $366 after combining the x -1.55x = -$24.80 after subtracting 391 from both sides Now solve for x (white) and 34-x the yellow.

OpenStudy (radar):

You don't have to be "good" at, all you need is to apply the rules for algebra that you learned in high school.

OpenStudy (radar):

To solve for x, divide both sides by -1.55

OpenStudy (anonymous):

x=16

OpenStudy (radar):

yes 16 white 34-16= 18 yellow

OpenStudy (anonymous):

thank you that helped i have a problem with percentages can u help

OpenStudy (radar):

go ahead and post it

OpenStudy (anonymous):

5900 is invested part of it @ 10% and part at 8% for a certain year the total yeild $534.00 how much was invested @ each rate?

OpenStudy (radar):

I think I have to think a little on this one. I'm thinking.

OpenStudy (anonymous):

ok i don't understand percantages thank you for helping me. i still have 3 graphs to do n i do not get graphing

OpenStudy (radar):

O. K. Lets do it this way Let x equal the part that was invested at 10% that being the case then we can say: 5900-x equals the part invested at 8 % We can say: .1x +.08(5900-x)=$534.00 You can take it from there.

OpenStudy (radar):

Graphs you will have to do your self, just plot the data.